# EMERGE: Integrating RAG for Improved Multimodal EHR Predictive Modeling

### v1

@@ 4.2.2 Retrieve Information from External KG
$\bm{h}_{n}$ $={\rm TextEncoder}(n),n\in Nodes$ (8)
$\bm{h}_{e}$ $={\rm TextEncoder}(e),e\in E$ (9)
@@
where $n$ and $e$ symbolize disease entities from $Nodes$ and the extracted entity set. Respectively, ${h}_{n}$ and ${h}_{e}$ denote their corresponding embeddings.
@@
When match relative nodes, we take current entity $e$ (including abnormal features and potential disease names) as the query. Then we compute the similarities between $e$ and each node in KG. The metric used for these calculation is cosine similarity, as shown below:
$\theta^{n}_{e}=\frac{{\bm{h}_{n}}\cdot{\bm{h}_{e}}}{\|{\bm{h}_{n}}\|\|{\bm{h}_{e}}\|}$ (10)
where $\bm{h}_{e}$ and $\bm{h}_{n}$ denote the embeddings of the entity $e$ and the node.
@@
In our approach, we establish a threshold to gauge the requisite similarity between two embeddings. We focus on nodes that surpass this threshold, ensuring that only the most relevant matches are considered:
$f(e,Nodes)=\begin{cases}\{n\in Nodes:\theta^{n}_{e}=\max_{n^{\prime}\in Nodes}\theta^{n^{\prime}}_{e}\}&\text{if }\max_{n\in Nodes}\theta^{n}_{e}>\eta\\ \emptyset&\text{otherwise}\end{cases}$ (11)
@@
where $\eta$ is the threshold for similarity, and $f(e,Nodes)$ denotes the set of nodes that we exclusively accept as matches for the entity $e$.
@@
Subsequently, we can obtain the definitions and descriptions within the disease entities, each represented as a node of the graph. Furthermore, relationships between diseases, encapsulated within triples, act as the edges of the graph. These pieces of information elaborate on the severity of the diseases, the harm they pose to the human body, and their interconnections from various perspectives. They further clarify the entity information from the original notes, thereby enhancing LLM’s understanding of the patient’s health condition.
4.2.3 Summarize and Encode KG Knowledge
@@
Drawing from the entities extracted from time-series and clinical notes, along with supplementary information about them, we have compiled extensive details about the patient’s medical condition. However, this content contains too many tokens for conventional language model inputs (such as BERT). As a countermeasure, we utilize retrieval-augmented generation to condense the aforementioned details, thereby attaining a concise representation of the patient’s health status.
@@
The prompt template, as illustrated in Figure 5, begins by defining a role and instructions to guild the generation by the LLM. Subsequently, we enumerate all abnormal features derived from the time-series data, and disease names extracted from clinical notes, which reflect the patient’s health threats. To enhance the comprehension, we integrate retrieved disease definitions and descriptions, along with the relationships sampled from KG to form a comprehensive supplementary resource. Based on this augmented information, the LLM compiles a summary of the patient’s health status.
@@
Finally, we employ a language model, denoted as TextEncoder, to encode the retrieval knowledge from external KG as below:
$\bm{h}_{RAG}={\rm TextEncoder}(\bm{x}_{RAG})$ (12)
@@
Figure 5: Prompt template for summary generation.
4.3 Multimodal Fusion Network
Figure 6: Fusion module. It combines multimodal embeddings with attention mechanism into a fused representation.
@@ 4.3 Multimodal Fusion Network
$\bm{h}_{Text}={\rm TextFusion}({\rm Concat}\left[\bm{h}_{Note},\bm{h}_{RAG}\right])$ (13)
To better integrate information from different modalities, we proposed an attention-based fusion network primarily consisting of cross-attention layers. First, the $Query$ vector is computed from the hidden representation of the other modality, while the $Key$ and $Value$ vectors are computed from the hidden representations of the current modality:
${Q}_{Text}=\bm{W}_{q}\cdot{h}_{Text}$ $,\ \ {Q}_{TS}=\bm{W}_{q}\cdot{h}_{TS}$ (14)
@@
$\bm{z}_{Text}$ $={\rm softmax}(\frac{{Q}_{TS}\bm{{K}}_{Text}^{\top}}{\sqrt{d_{k}}})\cdot{V}_{Text}$ (15)
@@
As a result, the outputs of the two cross-attention modules have carried information from both modalities. We further concatenate them and use MLP layer to obtain the fused information.
$\bm{z}$ $={\rm MLP}({\rm Concat}\left[\bm{z}_{TS},\bm{z}_{Text}\right])$ (16)
@@
$\mathcal{L}(\hat{y},y)=-\frac{1}{N}\sum_{i=1}^{N}(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i}))$ (18)
where $N$ is the number of patients within one batch, $\hat{y}\in[0,1]$ is the predicted probability and $y$ is the ground truth.
@@
5 Experimental Setups
5.1 Experimented Datasets and Utilized KG
Sourced from the EHRs of the Beth Israel Deaconess Medical Center, MIMIC-III and MIMIC-IV dataset is extensive and widely used in healthcare research. We adhere to the established EHR benchmark pipeline [40, 41] for preprocessing time-series data. 17 lab test features (include categorical features) and 2 demographic features (age and gender) are extracted. To minimize missing data, we consolidate every consecutive 12-hour segment into a single record for each patient, focusing on the first 48 records. And we follow Clinical-LongFormer[42]’s approach to extract and preprocess clinical notes, which includes minimal but essential steps: removing all de-identification placeholders to protect Protected Health Information (PHI), replacing non-alphanumeric characters and punctuation marks, converting all letters to lowercase for consistency, and stripping extra white spaces.
@@ 5.1 Experimented Datasets and Utilized KG
Table 2: Statistics of datasets after preprocessing. The number and proportion for labels are the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
@@
The external knowledge base we utilized is PrimeKG [25], which integrates 20 high-quality resources to describe 17,080 diseases with 4,050,249 relationships representing ten major biological scales, including disease-associated entities. Futhermore, PrimeKG extracts textual features of disease nodes containing information about disease prevalence, symptoms, etiology, risk factors, epidemiology, clinical descriptions, management and treatment, complications, prevention, and when to seek medical attention, which are highly relevant to the clinical prediction tasks.
@@
When prompting the LLM to generate the summary, 4 patients in the MIMIC-III dataset were not successfully generated due to DeepSeek-v2 [43]’s strict content censor policy, which flagged “Content Exists Risk.” We replaced these with “None”.
5.2 Evaluation Metrics
  • AUROC: This metric is our primary consideration in binary classification tasks due to its widespread use in clinical settings and its effectiveness in handling imbalanced datasets [44].
  • AUPRC: The AUPRC is particularly useful for evaluating performance in datasets with a significant imbalance between classes [45].
  • min(+P, Se): This composite metric represents the minimum value between precision (+P) and sensitivity (Se), providing a balanced measure of model performance [46].
All these three metrics are the higher the better.
5.3 Baseline Models
5.3.1 EHR Prediction Models
We include multimodal EHR baseline models (MPIM [8], UMM [36], MedGTX [35], VecoCare[9], M3Care [6]) and approaches that incorporating external knowledge from KG (GRAM [10], KAME [11], CGL [38], KerPrint [16], MedPath [14], MedRetriever [37]), and LLM facilitated model GraphCare [22] as our baselines.
  • MPIM [8] (ICML-2023) models the irregularity of time intervals in time-series EHR data and clinical notes via gating mechanism and apply interleaved attention mechanism for modality fusion.
  • UMM [36] (MLHC-2023) introduces Unified Multi-modal Set Embedding (UMSE) and Modality-Aware Attention (MAA) with Skip Bottleneck (SB) to handle challenges in data embedding and missing modalities in multi-modal EHR.
  • MedGTX [35] (CHIL-2022) introduces a pre-trained model for joint multi-modal representation learning of structured and textual EHR data by interpreting structured EHR data as a graph and employing a graph-text multi-modal learning framework.
  • VecoCare [9] (IJCAI-2023) addresses the challenges of synthesizing information from structured and unstructured EHR data using a Gromov-Wasserstein Distance-based contrastive learning and an adaptive masked language model.
  • M3Care [6] (KDD-2022) proposes an end-to-end model to handle missing modalities in multimodal healthcare data by imputing task-related information in the latent space using a task-guided modality-adaptive similarity metric.
  • GRAM [10] (KDD-2017) enhances EHRs with medical ontologies, using an attention mechanism to represent medical concepts by their ancestors in the ontology for improved predictive performance and interpretability.
  • KAME [11] (CIKM-2018) employs a knowledge attention mechanism to learn embeddings for nodes in a knowledge graph, improving the accuracy and robustness of health information prediction while providing interpretable disease representations.
  • CGL [38] (IJCAI-2021) proposes a collaborative graph learning model to explore patient-disease interactions and incorporate medical domain knowledge, integrating unstructured text data for accurate and explainable health event predictions.
  • KerPrint [16] (AAAI-2023) offers retrospective and prospective interpretations of diagnosis predictions through a time-aware KG attention method and an element-wise attention method for selecting candidate global knowledge, enhancing interpretability.
  • MedPath [14] (WWW-2021) employs graph neural networks to capture and integrate high-order connections from knowledge graphs into input representations, thereby enhancing the relevance and utility of external knowledge.
  • MedRetriever [37] (CIKM-2021) enhances health risk prediction and interpretability by leveraging unstructured medical text from authoritative sources. It combines EHR embeddings with features from target disease documents to retrieve relevant text segments.
  • GraphCare [22] (ICLR-2024) generates patient-specific knowledge graphs from external KGs and LLMs to improve healthcare predictions with a Bi-attention Augmented (BAT) GNN, demonstrating significant improvements in various healthcare prediction tasks.
5.3.2 Multimodal Fusion Methods
To examine the effectiveness of our fusion network, we consider fusion methods: Add [47], Concat [48, 49], Tensor Fusion (TF) [50], and MAG [51, 52].
  • Add [47]: Simply performs element-wise addition of features from different modalities to integrate information.
  • Concat [48, 49]: The representations are concated to do predictions.
  • Tensor Fusion (TF) [50]: It integrates information from multiple sources or modalities by creating a multimodal tensor representation that captures the interactions between these modalities.
  • MAG [51, 52]: It dynamically fuse information from different modalities by adaptively weighting and integrating the features from each modality based on their relevance and contribution to the task at hand.
5.4 Implementation Details
5.4.1 Hardware and Software Configuration
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 12.4. The server’s system memory (RAM) size is 128GB. We implement the model in Python 3.8.19, PyTorch 2.2.2 [53], PyTorch Lightning 2.2.4 [54], and pyehr [41].
5.4.2 Model Training and Hyperparameters
AdamW [55] is employed with a batch size of 256 patients. All models are trained for 100 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${32,64,128,256}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for EMERGE is: 128 hidden dimensions, 0.001 learning rate. The dropout rate is set to $0.25$. Performance is reported in the form of mean±std by applying bootstrapping on all test set samples 10 times for the MIMIC-III and MIMIC-IV datasets, following practices in AICare [56]. The threshold $\epsilon$ for identifying anomalies in time-series data is set as 2 (z-score value=2). The threshold $\eta$ for matching entities in KG is set as 0.6 for MIMIC-III and 0.7 for MIMIC-IV.
5.4.3 Utilized (Large) Language Models
EMERGE utilizes both Language Models (LMs) and Large Language Models (LLMs) in the pipeline. For LMs, we use the frozen-parameter pretrained Clinical-LongFormer [42]’s [CLS] token [29] for extracting textual embeddings and BGE-M3 [57] as the text embedding model to compute entity embeddings. For LLMs, we deploy an offline Qwen-7B [58] to extract entities from clinical notes and call the DeepSeek-V2 Chat [43] API to generate summaries.
  • Clinical-LongFormer [42]: Pretrained on MIMIC-III clinical notes, it is a domain-enriched language model designed to handle long clinical texts by extending the maximum input sequence length from 512 to 4096 tokens. Adopting it to extract MIMIC text embeddings follows Zhang et al. [8].
  • BGE-M3 [57]: Targets multi-language and cross-language text vectorization using a vast, diverse dataset. It is specifically designed for feature extraction and retrieval, accepting up to 8192 tokens with a hidden dimension of 1024.
  • Qwen 1.5-7B Chat [58]: A large language model pretrained on 3TB of data, supporting extensive task adaptability. It can handle contexts up to 8192 tokens, making it capable of processing patients’ clinical notes records.
  • DeepSeek-V2 Chat [43]: DeepSeek-V2 is a Mixture-of-Experts language model with 236B parameters, of which only 21B are activated per token, featuring architectures like Multi-head Latent Attention and DeepSeekMoE for efficient inference and economical training.
6 Experimental Results and Analysis
6.1 Experimental Results
The performance of our EMERGE framework on in-hospital mortality and 30-day readmission prediction tasks on the MIMIC-III dataset is summarized in Table 3. EMERGE consistently outperforms the baseline models, indicating its superior practical applicability in real-world clinical settings.
@@ 6.1 Experimental Results
Table 3: In-hospital mortality and 30-day readmission prediction results on the MIMIC-III and MIMIC-IV datasets. Bold indicates the best performance. All metrics are multiplied by 100 for readability purposes.
@@
6.2 Ablation Studies
6.2.1 Comparing Different Modality Fusion Strategies

### v2

@@ 4.2.2. Retrieve Information from External KG
(8) $\bm{h}_{n}$ $={\rm TextEncoder}(n),n\in Nodes$
(9) $\bm{h}_{e}$ $={\rm TextEncoder}(e),e\in E$
@@
where $n$ and $e$ symbolize disease entities from $Nodes$ and the extracted entity set, respectively. ${h}_{n}$ and ${h}_{e}$ denote their corresponding embeddings.
@@
When matching relative nodes, we take the current entity $e$ (including abnormal features and potential disease names) as the query. Then, we compute the similarities between $e$ and each node in the KG. The metric used for these calculations is cosine similarity:
(10) $\theta^{n}_{e}=\frac{{\bm{h}_{n}}\cdot{\bm{h}_{e}}}{\|{\bm{h}_{n}}\|\|{\bm{h}_{e}}\|}$
where $\bm{h}_{e}$ and $\bm{h}_{n}$ are embeddings of the entity $e$ and the node.
@@
We establish a threshold to gauge the requisite similarity between two embeddings. We focus on nodes that surpass this threshold, ensuring that only the most relevant matches are considered:
(11) $f(e,\text{Nodes})=\begin{cases}\{\hat{n}\}&\text{if }\theta^{\hat{n}}_{e}>\eta,\\ \emptyset&\text{otherwise},\end{cases}$
@@
where $\hat{n}=\arg\max_{n\in\text{Nodes}}\theta^{n}_{e}$, $\eta$ is the threshold for similarity, and $f(e,\text{Nodes})$ denotes the set of nodes that we exclusively accept as matches for the entity $e$.
@@
Subsequently, we can obtain the definitions and descriptions within the disease entities, each represented as a node of the graph. Furthermore, relationships between diseases, encapsulated within triples, act as the edges of the graph. These pieces of information elaborate on the severity of the diseases, the harm they pose to the human body, and their interconnections from various perspectives. They further clarify the entity information from the original notes, thereby enhancing the LLM’s understanding of the patient’s health condition.
4.2.3. Summarize and Encode KG Knowledge
@@
Drawing from the entities extracted from time-series data and clinical notes, along with supplementary information about them, we have compiled extensive details about the patient’s medical condition. However, this content contains too many tokens for conventional language model inputs (such as BERT). As a countermeasure, we utilize retrieval-augmented generation to condense the aforementioned details, thereby attaining a concise representation of the patient’s health status.
@@
The prompt template, as illustrated in Figure 5, begins by defining a role and instructions to guide the generation by the LLM. Subsequently, we enumerate all abnormal features derived from the time-series data, and disease names extracted from clinical notes, which reflect the patient’s health threats. To enhance comprehension, we integrate retrieved disease definitions and descriptions, along with the relationships sampled from the KG to form a comprehensive supplementary resource. Based on this augmented information, the LLM compiles a summary of the patient’s health status.
@@
Finally, we employ a language model, denoted as TextEncoder, to encode the retrieved knowledge from the external KG as below:
(12) $\bm{h}_{RAG}={\rm TextEncoder}(\bm{x}_{RAG})$
@@
Figure 5. Prompt template for summary generation.
4.3. Multimodal Fusion Network
Figure 6. Fusion module. It combines multimodal embeddings with attention mechanism into a fused representation.
@@ 4.3. Multimodal Fusion Network
(13) $\bm{h}_{Text}={\rm TextFusion}({\rm Concat}\left[\bm{h}_{Note},\bm{h}_{RAG}\right])$
To better integrate information from different modalities, we propose an attention-based fusion network primarily consisting of cross-attention layers. First, the $Query$ vector is computed from the hidden representation of the other modality, while the $Key$ and $Value$ vectors are computed from the hidden representations of the current modality:
(14) ${Q}_{Text}=\bm{W}_{q}\cdot{h}_{Text}$ $,\ \ {Q}_{TS}=\bm{W}_{q}\cdot{h}_{TS}$
@@
(15) $\bm{z}_{Text}$ $={\rm softmax}(\frac{{Q}_{TS}\bm{{K}}_{Text}^{\top}}{\sqrt{d_{k}}})\cdot{V}_{Text}$
@@
As a result, the outputs of the two cross-attention modules have carried information from both modalities. We further concatenate them and use an MLP layer to obtain the fused information:
(16) $\bm{z}$ $={\rm MLP}({\rm Concat}\left[\bm{z}_{TS},\bm{z}_{Text}\right])$
@@
(18) $\mathcal{L}(\hat{y},y)=-\frac{1}{N}\sum_{i=1}^{N}(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i}))$
where $N$ is the number of patients within one batch, $\hat{y}\in[0,1]$ is the predicted probability, and $y$ is the ground truth.
@@
5. Experimental Setups
5.1. Experimented Datasets and Utilized KG
Sourced from the EHRs of the Beth Israel Deaconess Medical Center, MIMIC-III and MIMIC-IV dataset is extensive and widely used in healthcare research. We adhere to the established EHR benchmark pipeline (Gao et al., 2024; Zhu et al., 2024b) for preprocessing time-series data. 17 lab test features (include categorical features) and 2 demographic features (age and gender) are extracted. To minimize missing data, we consolidate every consecutive 12-hour segment into a single record for each patient, focusing on the first 48 records. And we follow Clinical-LongFormer(Li et al., 2023)’s approach to extract and preprocess clinical notes, which includes minimal but essential steps: removing all de-identification placeholders to protect Protected Health Information (PHI), replacing non-alphanumeric characters and punctuation marks, converting all letters to lowercase for consistency, and stripping extra white spaces.
@@ 5.1. Experimented Datasets and Utilized KG
Table 1. Statistics of datasets after preprocessing. The number and proportion for labels are the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
@@
The external knowledge base we utilized is PrimeKG (Chandak et al., 2023), which integrates 20 high-quality resources to describe 17,080 diseases with 4,050,249 relationships representing ten major biological scales, including disease-associated entities. Futhermore, PrimeKG extracts textual features of disease nodes containing information about disease prevalence, symptoms, etiology, risk factors, epidemiology, clinical descriptions, management and treatment, complications, prevention, and when to seek medical attention, which are highly relevant to the clinical prediction tasks.
@@
When prompting the LLM to generate the summary, 4 patients in the MIMIC-III dataset were not successfully generated due to DeepSeek-v2 (DeepSeek-AI, 2024)’s strict content censor policy, which flagged “Content Exists Risk.” We replaced these with “None”.
@@
Table 2. In-hospital mortality and 30-day readmission prediction results on the MIMIC-III and MIMIC-IV datasets. Bold indicates the best performance. All metrics are multiplied by 100 for readability purposes.
@@
5.2. Evaluation Metrics
  • AUROC: This metric is our primary consideration in binary classification tasks due to its widespread use in clinical settings and its effectiveness in handling imbalanced datasets (McDermott et al., 2024).
  • AUPRC: The AUPRC is particularly useful for evaluating performance in datasets with a significant imbalance between classes (Kim and Hwang, 2022).
  • min(+P, Se): This composite metric represents the minimum value between precision (+P) and sensitivity (Se), providing a balanced measure of model performance (Ma et al., 2022).
All these three metrics are the higher the better.
5.3. Baseline Models
5.3.1. EHR Prediction Models
We include multimodal EHR baseline models (MPIM (Zhang et al., 2023a), UMM (Lee et al., 2023), MedGTX (Park et al., 2022), VecoCare(Xu et al., 2023), M3Care (Zhang et al., 2022)) and approaches that incorporating external knowledge from KG (GRAM (Choi et al., 2017), KAME (Ma et al., 2018), CGL (Lu et al., 2021), KerPrint (Yang et al., 2023), MedPath (Ye et al., 2021a), MedRetriever (Ye et al., 2021b)), and LLM facilitated model GraphCare (Jiang et al., 2024) as our baselines.
5.3.2. Multimodal Fusion Methods
To examine the effectiveness of our fusion network, we consider fusion methods: Add (Wu and Han, 2018), Concat (Khadanga et al., 2019; Deznabi et al., 2021), Tensor Fusion (TF) (Zadeh et al., 2017), and MAG (Rahman et al., 2020; Yang and Wu, 2021).
5.4. Implementation Details
5.4.1. Hardware and Software Configuration
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 12.4. The server’s system memory (RAM) size is 128GB. We implement the model in Python 3.11, PyTorch 2.2.2 (Paszke et al., 2019), PyTorch Lightning 2.2.4 (Falcon, 2019), and pyehr (Zhu et al., 2024b).
5.4.2. Model Training and Hyperparameters
AdamW (Loshchilov and Hutter, 2017) is employed with a batch size of 256 patients. All models are trained for 100 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${32,64,128,256}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for EMERGE is: 128 hidden dimensions, 0.001 learning rate. The dropout rate is set to $0.25$. Performance is reported in the form of mean±std by applying bootstrapping on all test set samples 10 times for the MIMIC-III and MIMIC-IV datasets, following practices in Ma et al. (2023). The threshold $\epsilon$ for identifying anomalies in time-series data is set as 2 (z-score value=2). The threshold $\eta$ for matching entities in KG is set as 0.6 for MIMIC-III and 0.7 for MIMIC-IV.
5.4.3. Utilized (Large) Language Models
EMERGE utilizes both Language Models (LMs) and Large Language Models (LLMs) in the pipeline. For LMs, we use the frozen-parameter pretrained Clinical-LongFormer (Li et al., 2023)’s [CLS] token (Devlin et al., 2018) for extracting textual embeddings and BGE-M3 (Chen and Xiao, 2024) as the text embedding model to compute entity embeddings. For LLMs, we deploy an offline Qwen-7B (Bai et al., 2023a) to extract entities from clinical notes and call the DeepSeek-V2 Chat (DeepSeek-AI, 2024) API to generate summaries.
6. Experimental Results and Analysis
6.1. Experimental Results
The performance of our EMERGE framework on in-hospital mortality and 30-day readmission prediction tasks on the MIMIC-III and MIMIC-IV datasets is summarized in Table 2. EMERGE consistently outperforms the baseline models, indicating its superior practical applicability in real-world clinical settings.
6.2. Ablation Studies
6.2.1. Comparing Different Modality Fusion Strategies